# Machine-learning correction for the calorimeter saturation of cosmic-rays ions with the Dark Matter Particle Explorer: towards the PeV scale

### v1

@@ 5 Conclusion and discussion
An ML method has been developed to correct the deposited energies in a calorimeter, in the presence of saturation. The issue at hand provokes the partial loss of the original energy deposits, due to the inability of the readout electronics to process signals above their maximum capacity. The use of a CNN makes it possible to analyse the BGO image with saturation and recover the original total deposited energy.
Prior to this work, two other corrections were developed to tackle the saturation in DAMPE calorimeter. The work in [12] provides an analytical method to infer the energy lost in saturation using the information from the adjacent bars; the main limitation of this approach is the inability of application where several contiguous saturated bars are found. On the other hand, the work in [13] uses an ML technique to achieve a higher precision; however, the final correction needs the training of two separate CNNs to reconstruct the single energy deposits lost in the middle layers and in the last one of the calorimeter. Also, two specialised trainings are performed for protons and helium, and a bias correction is needed for other ions. Overall, both precedent works succeed in their aim, but they require specific tunings for different ions and start loosing accuracy above $\sim$500 TeV of incident energy, posing issues for the flux measurements at the PeV scale. In this work, the focus has been moved to the prediction of an overall correction factor for the total deposited energy, instead of reconstructing the single pixels of the BGO image. This novel approach facilitates the model learning, achieving an improved precision. In addition, the introduction of different ion species and higher incident energies in the model training results in an improved accuracy, up to several PeVs. Both the improvements in maximum energy and generalisation supersede previous techniques, and can significantly help in probing the flux of CR ions at the PeV scale, where the knee region—i.e. a change in the all-particle spectral index, is expected to be. On the other hand, highly saturated events with more than 90% of the original deposit lost still require a deeper understanding to eliminate the residual bias.
@@ Appendix A Events re-weighting
MC simulations generate events with a spectrum $\propto E_{\mathrm{inc}}^{-1}$, with $E_{\mathrm{inc}}$ being the incident energy of the generated particle. The used spectral index is not compatible with the observations, that are more compatible with a spectrum $\propto E_{\mathrm{inc}}^{-2.7}$. To re-weight the events to a new spectral index $\gamma$, the following weights can be applied:
@@
with $N_{\mathrm{tot}}$ the total number of events and $M_{\mathrm{samples}}$ the number of samples covering energy ranges which $E_{\mathrm{inc}}$ belongs to. Figure 9 shows the distributions of proton, carbon and iron events used in this work before and after re-weighting to a $E^{-2.7}$ spectrum.
Figure 9: distribution of the primary energy, before and after re-weighting the entries to a simil-CR spectrum. The events are re-weighted so that the flux $\Phi\propto E_{\mathrm{inc}}^{-2.7}$, with $E_{\mathrm{inc}}$ being the incident energy. Three plots are shown here for the protons (top left), carbon (top right) and iron events (bottom). The distributions consider the full simulated samples, after event selection.
@@ Appendix B Fraction of saturated events
Generally, BGO bars of DAMPE calorimeter are more subject to saturation as the incident energy increases. Figure 10 shows the fraction of saturated events found in MC simulations of protons, carbon and iron, in function of the incident energy.
Figure 10: Fraction of saturated events found in the MC simulations of protons, carbon and iron nuclei crossing DAMPE instrument. The fraction is plotted in function of the true incident energy, and in Figure 9 it is possible to see the original event distributions with the total counts.

### v2

@@ 5 Conclusion and discussion
An ML method is developed to correct the deposited energies in the DAMPE calorimeter, in the presence of saturation. The issue at hand provokes the partial loss of the original energy deposits, due to the inability of the readout electronics to process signals above their maximum capacity. The use of a CNN makes it possible to analyse the BGO image with saturation and recover the original total deposited energy.
Prior to this work, two other corrections were developed to tackle the saturation in DAMPE calorimeter. The work in [22] provides an analytical method to infer the energy lost in saturation, using the information from the adjacent bars; the main limitation of this approach is the inability of application where several contiguous saturated bars are found. On the other hand, the work in [18] uses an ML technique to achieve a higher precision, but the final correction relies on two different CNNs to recover the energy lost respectively in the middle layers or in the last one. The two models also need specific trainings for different ions. Overall, both precedent works succeed in their aim, but they require specific tunings for different ions and start losing accuracy above $\sim$500 TeV of incident energy, posing issues for the flux measurements at the PeV scale. In this work, the focus moved to the prediction of an overall correction factor for the total deposited energy, instead of reconstructing the single pixels of the BGO image. This novel approach facilitates the model learning, achieving an improved precision. In addition, the introduction of different ion species and higher incident energies in the model training results in an improved accuracy, up to several PeVs. Both the improvements in maximum energy and generalisation supersede previous techniques, and can significantly help in probing the flux of CR ions at the PeV scale, where the knee region—i.e., a change in the all-particle spectral index—is expected to be. On the other hand, highly saturated events, with more than 90% of the original deposit lost, still require a deeper understanding to eliminate the residual bias.
The comparison between pseudo-saturated MC and data events shows a consistent behavior of the saturation correction on simulated and real events.
Acknowledgements
The DAMPE mission was funded by the strategic priority science and technology projects in space science of Chinese Academy of Sciences. In Europe the activities and data analysis are supported by the Swiss National Science Foundation (SNSF), Switzerland (grant No. 204253), the National Institute for Nuclear Physics (INFN), Italy, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, and the Swiss State Secretariat for Education, Research and Innovation (SERI).
@@ Appendix A Events re-weighting
MC simulations generate events with a spectrum $\propto E_{\mathrm{inc}}^{-1}$, depending on the incident particle energy $E_{\mathrm{inc}}$. Experimental measurements suggest that a spectral index of -2.7 is more appropriate. To modify the spectral index of simulated events to $\gamma$, a weight $w$ can be associated with each entry based on the definition
@@
with $N_{\mathrm{tot}}$ the total number of events, and $M_{\mathrm{samples}}$ the number of different samples populating the same energy range to which $E_{\mathrm{inc}}$ belongs. Figure 12 shows the distributions of proton, carbon and iron events used in this work before and after re-weighting to an $E^{-2.7}$ spectrum.
Figure 12: distribution of the primary energy, before and after re-weighting the entries to a simil-CR spectrum, proportional to $E_{\mathrm{inc}}^{-2.7}$. Three different plots are shown for the proton (top left), carbon (top right) and iron events (bottom). The distributions consider the full simulated samples after event selection.
@@ Appendix B Fraction of saturated events
Generally, BGO bars of DAMPE calorimeter saturate more as the incident energy increases. Figure 13 shows the fraction of saturated events found in MC simulations of proton, carbon and iron, as a function of the incident energy.
Figure 13: relative fraction of saturated events in MC simulations of protons, carbon and iron nuclei crossing DAMPE instrument. The fraction is plotted as a function of the true incident energy. The absolute total number of events can be found in Figure 12.
Appendix C PSD charge calculation
The DAMPE PSD can provide up to 4 independent measurements of the traversing particle charge. From all the measured deposits in the scintillators of a sub-layer, the one closest to the particle track is selected for the average charge measurement. For tracking purposes, the ML algorithm described in [20] is employed. After locating the closest PSD hits to the candidate track, the deposits are corrected according to the traversed path in the scintillator material; hits with a path length shorter than 5 mm are discarded. Having obtained the vector of independent charge measurements, a top-to-bottom progressive average is calculated. A hit is permitted to differ from the partial average by 2. When an inconsistent hit is found, the particle has likely interacted inelastically with the detector material, therefore the observed hit and all subsequent ones are rejected, and the average is finalised.